Robert has just bought a new model rocket, and is trying to measure its flight characteristics. The rocket engine package claims that it will maintain a constant thrust of 11.1 N until the engine is used up. Robert launches the rocket on a windless day, so that it travels straight up, and uses his laser range-finder to meaure that the height of the rocket when the engine cuts off is 10.2 m. He also measures the rocket\'s peak height, which is 14.7 m. If the rocket has a mass of 0.613 kg, how much work is done by the drag force on the rocket during its ascent?
Answer #1

Sol::

Given data

constant thrust =11.1 N
height of the rocket = 10.2 m
rocket\'s peak height = 14.7 m
mass of = 0.613 kg

Total work of the rocket engine in lifting rocket = 11.1(10.2)
= 113.22 J

Max GPE of rocket = mgh
= 0.613(9.81)(14.7)
= 88.398 J

Lost work/energy in reaching rocket's max height = 113.22-88.398
= 24.822 J

The drag force does negative work (on rocket)

Hope this answer helps you
Please RATE my answer
Thank you